# Delving into Instance-Dependent Label Noise in Graph Data: A Comprehensive Study and Benchmark

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3711896.3737376conference: ACM SIGKDD; Aug. 03–07, 2025; Toronto, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Neural networksccs: Computing methodologies Semi-supervised learning settings
@@ Appendix B Details of Label Corruption Process
LLM-based label noise.   Section 2.4 discusses the generation of LLM-based label noise. Table 7 presents the prompts used in our study. We employ an LLM as a text classifier, providing explicit instructions to generate the most appropriate category along with a justification using GPT-4o-mini (OpenAI). To reduce sensitivity to variations in LLM responses, we refine the generated label noise by filtering out the simple case that can be correctly predicted without explicit reasoning. Specifically, we create LLM-based label noise twice: one (LLM-Naive) without requesting reasoning and the other (LLM-Reasoned) with requesting reasoning. Then, we only adopt the LLM-reasoned labels in cases where both the LLM-naive and LLM-reasoned outputs disagree with the ground-truth label. This ensures that the retained noisy labels reflect more complex cases by filtering out simple misclassifications that could be corrected without deeper analysis.
@@
Appendix C Additional Results for Robust Learning
Impact of different noise types and levels on GNN models.   Section 3.1 examines how different noise types affect the performance of GNN models. Table 8 along with Table 1 present node classification accuracy (%) for MLP, GCN, GAT, GraphSAGE, and GIN, providing a comprehensive performance comparison across models and datasets. This evaluation suggests insights into how each GNN architecture performs under various conditions, highlighting the impact of different noise types and dataset characteristics on classification accuracy. Furthermore, in Table 9 and Table 10, we evaluate the classification accuracy on LLM-Naive and LLM-Reasoned to analyze the impact of varying noise levels on model performance. Figure 10 presents the summarized results for GraphSAGE. Performance deteriorates as the noise rate increases, yet the overall trend remains consistent across different noise types, as discussed in Section 3.1. We exclude the Children dataset when calculating the average model accuracy for homophilic datasets, as its noise rate exceeds 0.5, making it an impractical scenario for pairwise noise.

### v2

@@ Abstract.
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3711896.3737376isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Neural networksccs: Computing methodologies Semi-supervised learning settings
@@ Appendix B Details of Label Corruption Process
LLM-based label noise.   Section 2.4 discusses the generation of LLM-based label noise. We employ an LLM as a text classifier, providing explicit instructions to generate the most appropriate category along with a justification using GPT-4o-mini (OpenAI). The prompt for Cora-ML is “You are a domain expert in Machine Learning. Your task is to predict the most appropriate topic label for a research paper. You must choose from the following topic labels only:
To reduce sensitivity to variations in LLM responses, we refine the generated label noise by filtering out the simple case that can be correctly predicted without explicit reasoning. Specifically, we create LLM-based label noise twice: one (LLM-Naive) without requesting reasoning and the other (LLM-Reasoned) with requesting reasoning. Then, we only adopt the LLM-reasoned labels in cases where both the LLM-naive and LLM-reasoned outputs disagree with the ground-truth label. This ensures that the retained noisy labels reflect more complex cases by filtering out simple misclassifications that could be corrected without deeper analysis.